OpenStudy (anonymous):

ΔABC is a right triangle. m ∠A is 2° less than three times m ∠B. Find m ∠A and m ∠B, assuming ∠A is an acute angle.

OpenStudy (anonymous):

what do you know about the sum of angles A and B?

OpenStudy (anonymous):

A right triangle, with c

OpenStudy (anonymous):

so angle A + angle B = how many degrees?

OpenStudy (math&ing001):

You have A+B=90 and A=3*B-2 so 3*B-2+B=90 ie B=23 and A=90-B=67